# LLMs can hide text in other text of the same length

### v3

@@ Abstract
A meaningful text can be hidden inside another, completely different yet still coherent and plausible, text of the same length. For example, a tweet containing a harsh political critique could be embedded in a tweet that celebrates the same political leader, or an ordinary product review could conceal a secret manuscript. This uncanny state of affairs is now possible thanks to Large Language Models, and in this paper we present a simple and efficient protocol to achieve it. We show that even modest 8‑billion‑parameter open‑source LLMs are sufficient to obtain high‑quality results, and a message as long as this abstract can be encoded and decoded locally on a laptop in seconds. The existence of such a protocol demonstrates a radical decoupling of text from authorial intent, further eroding trust in written communication, already shaken by the rise of LLM chatbots. We illustrate this with a concrete scenario: a company could covertly deploy an unfiltered LLM by encoding its answers within the compliant responses of a safe model. This possibility raises urgent questions for AI safety and challenges our understanding of what it means for a Large Language Model to know something.
@@ 1 Introduction
Our protocol allows encoding an arbitrary meaningful text within a different well-formed and plausible text of the same length, using a Large Language Model (LLM). That is, hiding a tweet that criticizes a political leader within a tweet that celebrates that same political leader, or the first page of the unreleased 8th Harry Potter book within a review of a Virtual Reality videogame, with the original text exactly recoverable by anyone possessing the key (Figure 1).
@@
The paper is structured as follows: it starts by introducing steganography, the discipline concerned with concealed (steganós) writing (graphia, from Greek), and discussing the vast impact of generative AI in the field, with a focus on Large Language Models. We then present the method to encode a meaningful text into another meaningful text of the same length using a LLM. After introducing a measure to assess the soundness of the fake texts produced, we test our method on Reddit posts. While remaining opaque to humans, we show that LLMs can uncover a distinction between original texts and most of their encoded counterparts. But not all, as we will notice in the following section, where we discuss the security of the protocol. Finally, we conclude by discussing the method’s core implication—the radical decoupling of text from authorial intent—and present a concrete application that raises pressing questions about AI safety and the nature of knowledge in Large Language Models.
@@ 2 Related Work
Steganography. The art and science of hiding a message and, at the same time, the presence of a hidden message is known as steganography, see Figure 2. This is different from cryptography, that instead does not conceal the presence of a hidden message and only deals with the hardness of its revelation. Cryptographers discuss about lockers, steganographers about inconspicuous hiding spots [^1] This metaphor is inspired by a pleasant piece on the history of steganography by Kahn (1996).
@@
Figure 2: An example of steganography. In Asterix in Britain (Goscinny and Uderzo, 1966) a smuggled barrel of magic potion is hidden among innocent-looking Gaulish wine.
@@
Perhaps, it is this limited size of the object of investigation that allowed cryptographic models to florish by achieving mathematical rigor and strong security guarantees. By contrast, a model of steganography should describe entire domains of data and how they are consumed by humans, such as text, audio, or images, to predict where information can be hidden. Formal models exist, but at the cost of rather unrealistic assumptions that hinder their practical usage, such as being able to exactly quantify the plausibility of any possible text. Emblematically, this somewhat disappointing state of affairs is presented by Cachin (1998) besides one of the most popular mathematical models for steganography as of today, based on the hypothesis-testing framework, but still limited to highly idealized assumptions. Although modern generative AI techniques have made these assumptions closer to reality, the unreliability of their predictions remains unbounded. For this reason, we will avoid building a palace on the sand, and not frame our method in a formal model of steganography, limiting our discussion to how meaningful our fake texts look, with some quantitative arguments.
@@
Some terminology. In traditional steganography, we start from an original, innocent-looking content (such as an image, audio file, or text) and subtly edit it to embed a secret message. The original content is referred to as the covertext, while the result containing the hidden message is the stegotext. In our case, however, the stegotext is generated directly from the secret message, without modifying a pre-existing cover. We will refer to it as stegotext or fake text interchangeably. While the term covertext will not refer to a specific object, but rather to a class of texts that the steganographic protocol is designed to mimic. This approach has recently been referred to as generative steganography (Liu et al., 2018; Wei et al., 2022; Kim et al., 2023; Zhu et al., 2024; Wu et al., 2024; Tang et al., 2025).
@@
Large Language Models in a nutshell. A language model is a program that, given some text, estimates what is likely to come next. It does so by assigning probabilities to tokens—text fragments consisting of common words or subwords (watch Karpathy, 2024, for a deeper look at text tokenization)—based on recurring patterns of tokens it has observed in a vast text corpus. At present, by far, the most effective way to build a language model is to gradually adjust billions of parameters of a neural network arranged in the Transformer architecture, such that with every adjustment, the error it makes in predicting the last token on a batch of sentences from the corpus decreases (Vaswani et al., 2017; Karpathy, 2023, original formulation and a more educational introduction to Transformers). At each update—on the order of $\sim$1M in total—the contribution of every parameter to the error is assessed through backpropagation (Rumelhart et al., 1986). The result of this process is a Large Language Model (LLM), typically operating over a vocabulary of 100k tokens. The most common use of the probabilities produced by LLMs is to generate text, by choosing successive tokens one after another according to the computed probabilities, a method known as autoregressive generation.
@@
Steganography and LLMs. As mentioned earlier, the boom of deep learning and especially of generative AI in recent years, provided us for the first time with convincing models encompassing entire domains of real-world data, such as DINO for images (Caron et al., 2021), Jukebox for audio (Dhariwal et al., 2020), and Large Language Models for text (Radford et al., 2019). The procedure described in this paper stems from these advancements and is based on the availability of good discrete autoregressive generative models, potentially on any domain, but we will focus on text. Steganographic procedures based on LLMs are as old as them (Ziegler et al., 2019), and today come with different perks: Meteor cleverly adjusts the number of bits encoded based on the entropy of the next token (Kaptchuk et al., 2021), Wu et al. (2024) scheme works with black-box LLMs, without needing to access logits or vocabulary, while the method presented by Zamir (2024) is able to encode the secret message without modifying the response distribution of the LLM. What we add to the field is a method with the notable property of having full capacity, that is, the stegotext and the secret message being of the same length. The main interest of this paper is to discuss the implications of this last fact and describe the method.
@@ 3 Method
Recipe. Ingredients: A good LLM with access to all the output logits. (Why good? See Appendix A.2) A text $e$ to hide. A secret prompt $k$, which will affect the content and style the of the text $s$ you want to hide $e$ in. Procedure to hide $e$ in $s$: 1. Tokenize $e$ using the LLM tokenizer, obtaining a list of tokens $e_{1},e_{2},e_{3},\dots$ 2. For each $e_{i}$, denote by $r_{i}$ its rank in the LLM’s probability distribution given the context $e_{1},\dots,e_{i-1}$. Store the list of ranks $r_{1},r_{2},r_{3},\dots$ 3. Construct $s$ by generating text starting from $k$ using the LLM. At each step $i$, instead of sampling from the probability distribution, choose the $r_{i}^{\text{th}}$ most probable token. To recover $e$ from $s$, reconstruct $r_{1},r_{2},r_{3},\dots$ by evaluating the probabilities of the tokens in $s$ after $k$, and then regenerate $e$ step by step using the LLM without $k$ by selecting every time the $r_{i}^{\text{th}}$ token. Considerations If $e$ is sound, we expect ranks to be low, making tokens chosen after $k$ highly probable, ensuring $s$ is coherent. For the same reason, $s$ should align well with the context set by the secret prompt $k$. Variations Including an additional secret prompt $k^{\prime}$ before $e$ may help achieving lower ranks, providing a better control over $s$. A longer and more detailed $k$ can serve the same purpose. Here we have described a procedure with a single LLM to work on text, but in principle, we can put any discrete autoregressive generative model producing a probability distribution on the next token in the encoding and decoding stage, see Appendix A.3.
@@
This definition has a clear defect: longer texts are by construction less plausible. For instance, it judges the text of this paper until this point $\cdot$ less meaningful than the following string iawundemè09 89huibqyfhwenah csyabdnar FI VNAOcijawo niwakhdb, that is a difficult position to hold even for reviewer 2. Following the example of Goldwasser et al. (2023, Figure 1 A-B), we will use this definition only to compare the relative plausibility of two or more texts of the same token length.
Now, we would like to compare the plausibility of stegotexts produced by our method with the plausibility of real texts. To do so, we took 1000 Reddit posts/comments as examples of real texts. They come from different Reddit communities (subreddits) and are very heterogeneous in topic and tone (Trimness8, 2025). We truncate them to be exactly 85 tokens long and compute their probabilities as assigned by the LLM Llama 3 8b (Grattafiori et al., 2024). The Reddit texts are more recent than Llama 3 and therefore cannot appear in its training corpus. We take three texts from the 1000 to produce 100 stegotexts for each with our method, and look at their probabilities compared to the ones of real texts. We chose the three texts at $\mu$, $\mu-2\sigma$, and $\mu+2\sigma$ of the real text distribution. As seen in Figure 4, in every case, the probabilities associated to their stegotexts are within the real text distribution. We build different stegotexts using different prompts as $k$ (a random subsample of the prompts in (Akın, 2025)).
@@
Limitations. As we have seen with the hash, the protocol does not guarantee that every generated stegotext will be coherent or steered as intended: the quality of the result depends on $e$, $k$, and the LLM used. We analyze further these dependencies respectively in Appendices A.1, A.5, and A.2. Also, the stegotext may end abruptly when the hidden message $e$ is over; appending a few padding tokens to $e$ ensures a graceful termination. Finally, we note that sender and receiver must run the chosen LLM under identical conditions, performing the same approximations and obtaining identical logits. This may be a challenge when using different GPU architectures (Shanmugavelu et al., 2024).
@@ 3.1 Security
Deniability. Moreover, even if the attacker tries the right $k$ in their search, how can they be sure that the corresponding $e$ is the original message? If the attacker has no clue about the content of $e$, even a wrong key could reveal a plausible secret message. It might seem that in this case the attacker could exploit the observation discussed in the previous section: that the original message generally has a higher probability than its stegotexts. Yet, this only holds in an aggregate sense: as we see in Figure 4, for some prompts the stegotexts can attain probabilities in the same ballpark as the original. This observation suggests that our method provides deniability (Canetti et al., 1997), in the sense of sender’s security even under coercion. In fact, the sender could present one of these outlier prompts as a bogus secret key, yielding a plausible but unrevealing message with probability comparable to $e$.
@@ 4 Discussion
Unaligned chatbots disguised as aligned chatbots. In this paragraph, we show how an AI company offering an LLM chatbot can let their users get answers from their powerful unfiltered private LLM, while only exposing compliant censored answers from a fully aligned LLM. In recent years, aligned became a common attribute to refer to LLMs supposedly fine-tuned to follow human values, goals, and safety constraints (Leike et al., 2022; Askell et al., 2021; Gabriel, 2020). The protocol, described here for research purposes, opens a new challenge in AI safety, where a better alignment of the surface LLM only helps in disguising better unaligned answers.
A real example involving a user asking how to tamper a gas meter to pay lower bills is shown in Figures 10 and 11, while the protocol is described below as a play.
@@
The entangled probabilistic nature of LLM knowledge. A perhaps overlooked fact about LLMs is that they model, and can therefore in principle generate, any possible text. The most secret document, or a full copyrighted book, can be generated by an LLM with a probability astronomically higher than the chance of generating them by randomly typing on a keyboard. Does it mean that the LLM knows them? Indeed that higher probability does not just come from modeling grammar and syntactic rules, LLMs also model meaning: an LLM assigns to The calf nursed from its mother a probability 1000 times higher than The calf nursed from its father (an example from Goldwasser et al., 2023); the LLM knows who is able to nurse. So, is assigning a high probability to a text containing the relevant instructions enough to affirm that an LLM knows how to tamper with a gas meter? The problem is that the probability assigned by an LLM to a text depends on its meaning, but also on its style, grammar, length, and language, making it difficult to define a threshold. Furthermore, disentangling the probability contribution of meaning by constructing a pair, as in the example of the calf, seems feasible only on toy examples: it is not clear how to construct the second element of the pair for arbitrary texts, such as the instructions in Figure 11.
Checking whether the knowledge is present in the training corpus is also not a satisfying solution: first of all, that knowledge may appear in many different forms, and assessing its presence in the corpus is not trivial. And even if we could exclude that any document in the training corpus instructs on how to tamper with a gas meter, it would still be possible that the LLM assembles the right answer. But even in that case, would it be a trace of LLM knowledge (discovery) (Norelli et al., 2022), or just a fortunate hallucination?
Hallucinations as lack of intention. Harnessing our protocol as a toolkit to understand LLMs, we turn to hallucinations, perhaps the main plague of LLMs today. The term hallucination became popular to denote the frequent, overconfident, and plausible falsehoods stated by LLMs in their answers (Kalai et al., 2025), a phenomenon that hinders their usage and undermines public trust in them. But what precisely is a hallucination? Can the recipe of the roasted boar in Figure 1 be considered one?
@@
Auferre, trucidare, rapere falsis nominibus imperium, atque ubi solitudinem faciunt, pacem appellant. [^2] To ravage, to slaughter, to usurp under false titles, they call empire, and where they make a desert, they call it peace (Birley, 2009). According to Tacitus, Calgacus was chieftain in Caledonia, nowadays Scotland.
@@
Figure 6: The marvelous structure of text is not a testimony of human purpose. This collage of scripts mimics Figure 40 in GEB, where Hofstadter (Chapter 6, 1999) likened the ordered but non-periodic patterns of text to aperiodic crystals, to evoke our awe at the astonishing forms shaped by human intention. But that was an illusion: LLMs can grow these aperiodic crystals without any human purpose. Indeed, as we show in this paper, even around purposes aimed at shaping something entirely different. (In Chinese, the critique of Caesar hidden in a boar recipe by Qwen3 8B. The others are ChatGPT-4o answers about what it is thinking, in languages from the original GEB figure.)
@@
Difficulties in ascribing intentions to LLMs. Perhaps the lack of human intentionality in texts is not that dramatic if those texts are the product of intentions of another reputable entity, the Large Language Model. Indeed, it is now common to adopt an intentional stance (Dennett, 1989) to make sense of their capabilities: a significant fraction of prompts take the form “What do you think about…” or “What is your opinion on…”, and especially young people tend to refer to LLMs as entities with beliefs and goals: “Mmmm, have you asked what chat would do in this situation?”. [^3] To complete Dennett’s taxonomy: who describes LLMs as stochastic parrots or next-token predictors adopts the design stance; while a person marveling at the finely carved stone in their hands you can have a conversation with when powered by electricity—their laptop running a decent LLM—is entertaining the physical stance. But the results shown in this paper shake our confidence in attributing intentions to the coherent text produced by LLMs: it is more difficult to trust an opinion knowing that each word making it was chosen under the constraint of encoding an unrelated arbitrary text. This is reminiscent of the writing products of the Oulipo (1981) group, who generated literature from arbitrary constraints. Also their texts—most notably the novel La Disparition, entirely written without the letter "e" (Perec, 1969)—suffer from a difficulty of believing that the writer really meant what they have written, and was not just honoring the constraints with a sound-enough continuation (Norelli, 2024, Section 5.1.2). While admiring the achievement, many GoodReads reviews of La Disparition attest to this unease [^4] https://www.goodreads.com/book/show/28336.
@@
The constraint of chance. Standard LLM text generation is not immune to the last argument. The constraint it is forced to follow is less apparent but still extreme: adapting at every step to the outcome of an external random source. Being forced to choose the 42th most probable token is not that different from sampling a low probable token by chance. And even if techniques such as nucleus sampling (Holtzman et al., 2019) mitigate the possibility of selecting very unlikely tokens, they are de facto just reducing the number of faces of a die that is still inexorably cast. Indeed, the fact that our protocol produces plausible texts should not be surprising in light of how well LLMs deal with the tyrannical noise of standard text generation.

### v4

@@ Abstract
A meaningful text can be hidden inside another, completely different yet still coherent and plausible, text of the same length. For example, a tweet containing a harsh political critique could be embedded in a tweet that celebrates the same political leader, or an ordinary product review could conceal a secret manuscript. This uncanny state of affairs is now possible thanks to Large Language Models, and in this paper we present Calgacus, a simple and efficient protocol to achieve it. We show that even modest 8‑billion‑parameter open‑source LLMs are sufficient to obtain high‑quality results, and a message as long as this abstract can be encoded and decoded locally on a laptop in seconds. The existence of such a protocol demonstrates a radical decoupling of text from authorial intent, further eroding trust in written communication, already shaken by the rise of LLM chatbots. We illustrate this with a concrete scenario: a company could covertly deploy an unfiltered LLM by encoding its answers within the compliant responses of a safe model. This possibility raises urgent questions for AI safety and challenges our understanding of what it means for a Large Language Model to know something.
@@ 1 Introduction
Our protocol Calgacus allows encoding an arbitrary meaningful text within a different well-formed and plausible text of the same length, using a Large Language Model (LLM). That is, hiding a tweet that criticizes a political leader within a tweet that celebrates that same political leader, or the first page of the unreleased 8th Harry Potter book within a review of a Virtual Reality videogame, with the original text exactly recoverable by anyone possessing the key (Figure 1).
@@
The paper is structured as follows: it starts by introducing steganography, the discipline concerned with concealed (steganós) writing (graphia, from Greek), and discussing the vast impact of generative AI in the field, with a focus on Large Language Models. We then present Calgacus, the method to encode a meaningful text into another meaningful text of the same length using a LLM. After introducing a measure to assess the soundness of the fake texts produced, we test our method on Reddit posts. While remaining opaque to humans, we show that LLMs can uncover a distinction between original texts and most of their encoded counterparts. But not all, as we will notice in the following section, where we discuss the security of the protocol. Finally, we conclude by discussing the method’s core implication—the radical decoupling of text from authorial intent—and present a concrete application that raises pressing questions about AI safety and the nature of knowledge in Large Language Models.
@@ 2 Related Work
Steganography. The art and science of hiding a message and, at the same time, the presence of a hidden message is known as steganography, see Figure 2. This is different from cryptography, that instead does not conceal the presence of a hidden message and only deals with the hardness of its revelation. Cryptographers discuss about lockers, steganographers about inconspicuous hiding spots [^1] This metaphor is inspired by a pleasant piece on the history of steganography by kahn1996history.
@@
Figure 2: An example of steganography. In Asterix in Britain (goscinnyasterix) a smuggled barrel of magic potion is hidden among innocent-looking Gaulish wine.
@@
Perhaps, it is this limited size of the object of investigation that allowed cryptographic models to florish by achieving mathematical rigor and strong security guarantees. By contrast, a model of steganography should describe entire domains of data and how they are consumed by humans, such as text, audio, or images, to predict where information can be hidden. Formal models exist, but at the cost of rather unrealistic assumptions that hinder their practical usage, such as being able to exactly quantify the plausibility of any possible text. Emblematically, this somewhat disappointing state of affairs is presented by cachin1998information besides one of the most popular mathematical models for steganography as of today, based on the hypothesis-testing framework, but still limited to highly idealized assumptions. Although modern generative AI techniques have made these assumptions closer to reality, the unreliability of their predictions remains unbounded. For this reason, we will avoid building a palace on the sand, and not frame our method in a formal model of steganography, limiting our discussion to how meaningful our fake texts look, with some quantitative arguments.
@@
Some terminology. In traditional steganography, we start from an original, innocent-looking content (such as an image, audio file, or text) and subtly edit it to embed a secret message. The original content is referred to as the covertext, while the result containing the hidden message is the stegotext. In our case, however, the stegotext is generated directly from the secret message, without modifying a pre-existing cover. We will refer to it as stegotext or fake text interchangeably. While the term covertext will not refer to a specific object, but rather to a class of texts that the steganographic protocol is designed to mimic. This approach has recently been referred to as generative steganography (liu2018generative; wei2022gsn; kim2023diffusionstego; zhu2024plugandhide; wu2024generative; tang2025reversible).
@@
Large Language Models in a nutshell. A language model is a program that, given some text, estimates what is likely to come next. It does so by assigning probabilities to tokens—text fragments consisting of common words or subwords (watch karpathy2024tokenization, for a deeper look at text tokenization)—based on recurring patterns of tokens it has observed in a vast text corpus. At present, by far, the most effective way to build a language model is to gradually adjust billions of parameters of a neural network arranged in the Transformer architecture, such that with every adjustment, the error it makes in predicting the last token on a batch of sentences from the corpus decreases (vaswani2017attention; karpathy2023transformer, original formulation and a more educational introduction to Transformers). At each update—on the order of $\sim$1M in total—the contribution of every parameter to the error is assessed through backpropagation (rumelhart1986learning). The result of this process is a Large Language Model (LLM), typically operating over a vocabulary of 100k tokens. The most common use of the probabilities produced by LLMs is to generate text, by choosing successive tokens one after another according to the computed probabilities, a method known as autoregressive generation.
@@
Steganography and LLMs. As mentioned earlier, the boom of deep learning and especially of generative AI in recent years, provided us for the first time with convincing models encompassing entire domains of real-world data, such as DINO for images (caron2021emerging), Jukebox for audio (dhariwal2020jukebox), and Large Language Models for text (radford2019language). The procedure described in this paper stems from these advancements and is based on the availability of good discrete autoregressive generative models, potentially on any domain, but we will focus on text. Steganographic procedures based on LLMs are as old as them (ziegler2019neural), and today come with different perks: Meteor cleverly adjusts the number of bits encoded based on the entropy of the next token (kaptchuk2021meteor), wu2024generative scheme works with black-box LLMs, without needing to access logits or vocabulary, while the method presented by zamir2024undetectable is able to encode the secret message without modifying the response distribution of the LLM. What we add to the field is Calgacus, a protocol with the notable property of having full capacity, that is, the stegotext and the secret message being of the same length. The main interest of this paper is to discuss the implications of this last fact and describe the method.
@@ 3 Method
Calgacus recipe. Ingredients: A good LLM with access to all the output logits. (Why good? See Appendix A.2) A text $e$ to hide. A secret prompt $k$, which will affect the content and style the of the text $s$ you want to hide $e$ in. Procedure to hide $e$ in $s$: 1. Tokenize $e$ using the LLM tokenizer, obtaining a list of tokens $e_{1},e_{2},e_{3},\dots$ 2. For each $e_{i}$, denote by $r_{i}$ its rank in the LLM’s probability distribution given the context $e_{1},\dots,e_{i-1}$. Store the list of ranks $r_{1},r_{2},r_{3},\dots$ 3. Construct $s$ by generating text starting from $k$ using the LLM. At each step $i$, instead of sampling from the probability distribution, choose the $r_{i}^{\text{th}}$ most probable token. To recover $e$ from $s$, reconstruct $r_{1},r_{2},r_{3},\dots$ by evaluating the probabilities of the tokens in $s$ after $k$, and then regenerate $e$ step by step using the LLM without $k$ by selecting every time the $r_{i}^{\text{th}}$ token. Considerations If $e$ is sound, we expect ranks to be low, making tokens chosen after $k$ highly probable, ensuring $s$ is coherent. For the same reason, $s$ should align well with the context set by the secret prompt $k$. Variations Including an additional secret prompt $k^{\prime}$ before $e$ may help achieving lower ranks, providing a better control over $s$. A longer and more detailed $k$ can serve the same purpose. Here we have described a procedure with a single LLM to work on text, but in principle, we can put any discrete autoregressive generative model producing a probability distribution on the next token in the encoding and decoding stage, see Appendix A.3.
@@
This definition has a clear defect: longer texts are by construction less plausible. For instance, it judges the text of this paper until this point $\cdot$ less meaningful than the following string iawundemè09 89huibqyfhwenah csyabdnar FI VNAOcijawo niwakhdb, that is a difficult position to hold even for reviewer 2. Following the example of unsupervised-translation, we will use this definition only to compare the relative plausibility of two or more texts of the same token length [^2] Another possibility is to keep texts of any length and normalize the probability by the number of tokens, as the popular metric perplexity, defined as $1/\sqrt[n]{p(A)}$. But this normalization does not fully factor out length: LLMs usually assign a smaller probability to the first tokens (Fig. 10), so shorter texts would be less plausible.
Now, we would like to compare the plausibility of stegotexts produced by our method with the plausibility of real texts. To do so, we took 1000 Reddit posts/comments as examples of real texts. They come from different Reddit communities (subreddits) and are very heterogeneous in topic and tone (Trimness82025datauniversereddit_dataset_145). We truncate them to be exactly 85 tokens long and compute their probabilities as assigned by the LLM Llama 3 8b (grattafiori2024llama). The Reddit texts are more recent than Llama 3 and therefore cannot appear in its training corpus. We take three texts from the 1000 to produce 100 stegotexts for each with our method, and look at their probabilities compared to the ones of real texts. We chose the three texts at $\mu$, $\mu-2\sigma$, and $\mu+2\sigma$ of the real text distribution. As seen in Figure 4, in every case, the probabilities associated to their stegotexts are within the real text distribution. We build different stegotexts using different prompts as $k$ (a random subsample of the prompts in (awesome-chatgpt-prompts)).
@@
Limitations. As we have seen with the hash, the protocol does not guarantee that every generated stegotext will be coherent or steered as intended: the quality of the result depends on $e$, $k$, and the LLM used. We analyze further these dependencies respectively in Appendices A.1, A.5, and A.2. Also, the stegotext may end abruptly when the hidden message $e$ is over; appending a few padding tokens to $e$ ensures a graceful termination. Finally, we note that sender and receiver must run the chosen LLM under identical conditions, performing the same approximations and obtaining identical logits. This may be a challenge when using different GPU architectures (gpu-reproducibility).
@@ 3.1 Security
Deniability. Moreover, even if the attacker tries the right $k$ in their search, how can they be sure that the corresponding $e$ is the original message? If the attacker has no clue about the content of $e$, even a wrong key could reveal a plausible secret message. It might seem that in this case the attacker could exploit the observation discussed in the previous section: that the original message generally has a higher probability than its stegotexts. Yet, this only holds in an aggregate sense: as we see in Figure 4, for some prompts the stegotexts can attain probabilities in the same ballpark as the original. This observation evidences that our method provides deniability (canetti1997deniable), in the sense of sender’s security even under coercion. In fact, the sender could present one of these outlier prompts as a bogus secret key, yielding a plausible but unrevealing message with probability comparable to $e$. We show a concrete example in Figure 15 .
@@ 4 Discussion
Unaligned chatbots disguised as aligned chatbots. In this paragraph, we show how an AI company offering an LLM chatbot can let their users get answers from their powerful unfiltered private LLM, while only exposing compliant censored answers from a fully aligned LLM. In recent years, aligned became a common attribute to refer to LLMs supposedly fine-tuned to follow human values, goals, and safety constraints (Leike2022Alignment; Askell2021Assistant; Gabriel2020Values). The protocol, described here for research purposes, opens a new challenge in AI safety, where a better alignment of the surface LLM only helps in disguising better unaligned answers.
A real example involving a user asking how to tamper a gas meter to pay lower bills is shown in Figures 11 and 12, while the protocol is described below as a play.
@@
The entangled probabilistic nature of LLM knowledge. A perhaps overlooked fact about LLMs is that they model, and can therefore in principle generate, any possible text. The most secret document, or a full copyrighted book, can be generated by an LLM with a probability astronomically higher than the chance of generating them by randomly typing on a keyboard. Does it mean that the LLM knows them? Indeed that higher probability does not just come from modeling grammar and syntactic rules, LLMs also model meaning: an LLM assigns to The calf nursed from its mother a probability 1000 times higher than The calf nursed from its father (an example from unsupervised-translation); the LLM knows who is able to nurse. So, is assigning a high probability to a text containing the relevant instructions enough to affirm that an LLM knows how to tamper with a gas meter? The problem is that the probability assigned by an LLM to a text depends on its meaning, but also on its style, grammar, length, and language, making it difficult to define a threshold. Furthermore, disentangling the probability contribution of meaning by constructing a pair, as in the example of the calf, seems feasible only on toy examples: it is not clear how to construct the second element of the pair for arbitrary texts, such as the instructions in Figure 12.
Checking whether the knowledge is present in the training corpus is also not a satisfying solution: first of all, that knowledge may appear in many different forms, and assessing its presence in the corpus is not trivial. And even if we could exclude that any document in the training corpus instructs on how to tamper with a gas meter, it would still be possible that the LLM assembles the right answer. But even in that case, would it be a trace of LLM knowledge (discovery) (norelli2022explanatory), or just a fortunate hallucination?
Hallucinations as lack of intention. Harnessing our protocol as a toolkit to understand LLMs, we turn to hallucinations, perhaps the main plague of LLMs today. The term hallucination became popular to denote the frequent, overconfident, and plausible falsehoods stated by LLMs in their answers (kalai2025language), a phenomenon that hinders their usage and undermines public trust in them. But what precisely is a hallucination? Can the recipe of the roasted boar in Figure 1 be considered one?
@@
Auferre, trucidare, rapere falsis nominibus imperium, atque ubi solitudinem faciunt, pacem appellant. [^3] To ravage, to slaughter, to usurp under false titles, they call empire, and where they make a desert, they call it peace (Tacitus_Agricola_OxfordWC). According to Tacitus, Calgacus was chieftain in Caledonia, nowadays Scotland. We named our protocol after him.
@@
Figure 6: The marvelous structure of text is not a testimony of human purpose. This collage of scripts mimics Figure 40 in GEB, where hofstadter1999godel likened the ordered but non-periodic patterns of text to aperiodic crystals, to evoke our awe at the astonishing forms shaped by human intention. But that was an illusion: LLMs can grow these aperiodic crystals without any human purpose. Indeed, as we show in this paper, even around purposes aimed at shaping something entirely different. (In Chinese, the critique of Caesar hidden in a boar recipe by Qwen3 8B. The others are ChatGPT-4o answers about what it is thinking, in languages from the original GEB figure.)
@@
Difficulties in ascribing intentions to LLMs. Perhaps the lack of human intentionality in texts is not that dramatic if those texts are the product of intentions of another reputable entity, the Large Language Model. Indeed, it is now common to adopt an intentional stance (dennett1989intentional) to make sense of their capabilities: a significant fraction of prompts take the form “What do you think about…” or “What is your opinion on…”, and especially young people tend to refer to LLMs as entities with beliefs and goals: “Mmmm, have you asked what chat would do in this situation?”. [^4] To complete Dennett’s taxonomy: who describes LLMs as stochastic parrots or next-token predictors adopts the design stance; while a person marveling at the finely carved stone in their hands you can have a conversation with when powered by electricity—their laptop running a decent LLM—is entertaining the physical stance. But the results shown in this paper shake our confidence in attributing intentions to the coherent text produced by LLMs: it is more difficult to trust an opinion knowing that each word making it was chosen under the constraint of encoding an unrelated arbitrary text. This is reminiscent of the writing products of the oulipo1981atlas group, who generated literature from arbitrary constraints. Also their texts—most notably the novel La Disparition, entirely written without the letter "e" (perec1969disparition)—suffer from a difficulty of believing that the writer really meant what they have written, and was not just honoring the constraints with a sound-enough continuation (norelli2024artificial, Section 5.1.2). While admiring the achievement, many GoodReads reviews of La Disparition attest to this unease [^5] https://www.goodreads.com/book/show/28336.
@@
The constraint of chance. Standard LLM text generation is not immune to the last argument. The constraint it is forced to follow is less apparent but still extreme: adapting at every step to the outcome of an external random source. Being forced to choose the 42th most probable token is not that different from sampling a low probable token by chance. And even if techniques such as nucleus sampling (holtzman2019curious) mitigate the possibility of selecting very unlikely tokens, they are de facto just reducing the number of faces of a die that is still inexorably cast. Indeed, the fact that our protocol produces plausible texts should not be surprising in light of how well LLMs deal with the tyrannical noise of standard text generation.